OpenStudy (ttp):

Two families paid a total of $44.00 to see a movie. Each adult ticket cost $8.00. If there were 3 adults and 4 children, how much did each child's ticket cost? $5.00 $5.50 $6.00 $6.50

OpenStudy (the_fizicx99):

In general they paid 44, 8 for each adult, theres 3 so 8*3 8*3 = 24 44-24 = 20 20/4 = 5 So $5 per child ticket

OpenStudy (anonymous):

total cost=44 adult=8 child=? 3(8)+4(c)=44 4c=20 c=5 therefore 5 children tickets